A compound has a molar absorptivity of 1.50 x 102 /(M• cm) . if 1.0 mL of a solution containing this compound is placed in 5.0 mL of water, this final solution produced and absorbance of 1.500. What is the concentration of this compound in the original solution? (cuvette = 1.0 cm)
Answer #1

Beer's law states that the absorbance of a solution is directly proportional to its concentration.

The equation for Beer's law is

A=ε l c

Where,

A is the absorbance

ε is the molar absorptivity

l is the path length

c is the concentration

Given

A=1.500

ε=1.50 x 102 /(M• cm)

l=1cm

c=?

Substituting the values in equation for Beer's law

A=ε l c

1.500 =1.50 x 102 /(M• cm) * 1cm * c

c=1.500 /[1.50 x 102 /(M• cm) * 1cm ]

=10-2M

This is the concentration of the final solution obtained by diluting 1mL of original solution with 5 mL water, therefore final volume is 6mL.

Dilution factor =Final volume(Vf) / Initial volume(Vi) = 6 mL/1mL = 6

Dilution factor = initial concentration of solution / final concentration of solution

6 =initial concentration of solution / 10-2M

initial concentration of solution =6* 10-2M = 0.06 M